Fame is 10 years older than EJ. Five years ago, EJ was two years more than one-fourth as old as Fame. What are their present ages.
Answer by ptaylor(2198)

Let x=EJ's age; x-5=EJ's age 5 years ago
Then x+10=Fame's age; x+10-5=x+5=fame's age 5 years ago
Now we are told the following:
x-5=(1/4)(x+5)+2 multiply each term by 4
4x-20=x+5+8 subtract x from and add 20 to each side
4x-20+20-x=5+8+20 collect like terms
3x=33 divide each side by 3
x=11----------------------------EJ's age
x+10=11+10=21----------------------Fame's age
CK
11-5=(1/4)(21-5)+2
6=4+2
6=6

Hope this helps---ptaylor